A 0.030-kg ice cube a "0^oC" is placed in an insulated box that contains a fixed quantity of steam at "100^oC" . When thermal equilibrium of this closed system is established, its temperature is found to be "23^oC" . Determine the original mass (in kg) of the steam at "100^oC."


1
Expert's answer
2020-08-20T09:48:16-0400

While the steam condenses to water, the ice melts and then water heats up:

Heat that steam gives out when condenses:


"Q_s=-m_sq."


Heat the steam gives out while cools to 23°C:


"Q_{ws}=cm_s(23-100)."


Heat the ice receives when melts:


"Q_i=m_i\nL."

Heat the water obtained from ice receives while heats:


"Q_{wi}=cm_i(23-0)."

The equilibrium of thermal energy looks like this:


"0=Q_s+Q_i+Q_{w1}+Q_{w2},\\
0=-m_sq+m_iL+cm_i(23-0)+cm_s(23-100),\\
m_s[q-c(23-100)]=m_i[L+c(23-0)].\\
\\\space\\m_s=m_i\frac{L+c(23-0)}{q-c(23-100)},\\\space\\
m_s=0.03\frac{333550+4200(23-0)}{2264700-4200(23-100)},\\\space\\
m_s=6.64\cdot10^{-3}\text{ kg}."


The mass of steam was 6.64 grams.